A 1m3 air tank is initially at 100 kPa and 17 degrees Celsius. It is to be filled by a compressor that delivers air at 700 kPa and 52 degrees Celsius. When the tank is filled it is allowed to reach thermal equilibrium and the tank is at 17 degrees Celsius and 600 kPa.

Determine the amount and direction of heat transfer between the tank and surroundings. (Hint: first find the initial mass of air in the tank, the final mass of air in the tank, and the mass of air that entered the tank. Remember the ideal gas law applies to this scenario.)
Answer #1

Initial conditions P T of tank = 100 kPa = 1 m3 = 15°c = 990K Initial Ру mass a R=0.287 m, = - RJ, 100x 1 0.289 X 290 *, -1-2Final conditions V₂ =Vi= ims P2 = 600kPa T2 = 17° c = 290k m = P2 V₂ = RT 2 600x 1 29080.287 2 = 7.2089 kg Mass of compresseTQ = dU = Of - Vi = 1501.03 - 1652.00 (do =-100.97 KT | heat flow from system to - indicates surroundings.